Apr 20, 2017 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Disable red LED string in dark room to avoid leaving it on in the future. (Adesh)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
  3. Action Items
  4. Activities during the down
  5. FCAL update
  6. BCAL update
    1. Energy Resolution update - George
  7. Calibrations
  8. Reconstruction and Simulation
  9. NIM articles
  10. Any other business


Attendees: Elton, Jon, Adesh, Mark (JLab); Matt (IU); Sean (NW); George (Athens); Will (CMU); Teagan, Ahmed? (UofR)

  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
    • Schedule two talks 30 min each. Will (Bcal), Adesh (Fcal)
  3. Action Items
    • Will review Fcal base replacement and red light in darkroom with technical staff immediately following collaboration meeting
  4. Activities during the down
    • Mark: Perhaps we can diagnose the Bcal problematic channels during the down. Need to think about how this can be accomplished
    • Elton: Any groups that need to do work on the detector need to define effort and schedule, because the detector might get shutdown for a period.
  5. FCAL update
    1. Calibrations (Adesh)
      • All gain and non-linear corrections have been completed and are in the ccdb database.
      • Attention turned to timing issues.
      • Peak at -2ns comes from low energy showers, but source is unknown.
      • FADC 4 ns timing changes have been investigated using LED pulser. Find that only about 1% of channels are affected. About 115 channels show this behavior, with about 54 channels producing more than 10 changes during the spring. Channels with timing shifts are broadly distributed over the calorimeter. Plan to ad a list into ccdb of shifts relative to an arbitrary reference run.
      • Mark: Bcal finds timing shifts in the downstream, i.e. in the first 8 channels of the ADC. Adesh will check the location of shifting channels in the Fcal.
      • Adesh hopes to have the shift corrections by the collaboration, but probably not by the next launch.
  6. BCAL update
    1. Timing calibration (Mark)
      • Shifts of 4 ns in the Bcal have been problematic and affect higher level calibrations.
      • New technique to address the problem compares upstream to downstream channels and correcting for the position of the hit using tracking.
      • Needs about 100k events from a file for the calibration.
      • There are still some remaining issues with putting and pulling constants from ccdb, but should be fixed shortly. Procedure is close, but not completed. Hope to be ready for a launch over the weekend.
      • This new technique does not rely on the TDCs and is applicable to all 4 layers.
    2. Bcal hadronic Efficiencies (Sean)
      • After various fixes to the code, Sean still sees some dips in the phi distribution of tracks.
      • Elton: The dip was reduced considerably after Simon made some fixes, but was still visible. There were no indications this affected the layer efficiencies. However, the tracks to require a matched shower, so the Bcal could be affecting the overall distribution.
      • Mark: Subtle effects of the Bcal timing could be having an impact.
      • Need to continue to monitor the phi distribution of tracks and track down any phi asymmetries.
    3. Energy Resolution update - George
      • Generated 1M gp -> p pi0 -> p e+e- g events with Geant4 and reaction selected via analysis cuts. Only 12500 events remain after analysis, i.e. about 1% reconstruction efficiency.
      • Elton: Note that Geant4 resolutions for forward tracks seems to be very poor compared to Geant3. Might wish to check the process with Geant3 for comparison.
      • Matt: concerned that with the small branching fraction and the small analysis yield, statistics might be very low.
      • Sean: Has a pi0 Dalitz decay procedure for Geant3, which he will pass on to George for reference.
  7. Calibrations
    1. Pi0 gain calibration update (Will)
      • Updated Log Entry 3469359.
      • Updated layer 3 calibrations. Added Justin's saturation correction for large pulses and checked that non-linear corrections were essentially unchanged.
      • Mark: Check to see if the new 2017 gain constants improve the resolution for the 2016 data. If so, one could consider using them for the 2016 reconstruction launch. Will: check this weekend.
      • Mark: On a future launch, we may wish to work with J/Psi group to obtain sample of high E electrons to study systematics of the saturation.
      • Elton: Has Tegan followed up on shadowing Will for the gain calibration? Tegan: Checked MC to see if one could use an exclusive channel for the gain calibrations. This effort did not bear much fruit so is not being pursued.
  8. Reconstruction and Simulation
  9. NIM articles
  10. Any other business